# Retrieval of fundamental material parameters of monolayer transition metal dichalcogenides from experimental exciton energies: An analytical approach

### v1

@@ Abstract
We propose a straightforward and highly accurate method for determining material parameters such as screening length, bandgap energy, exciton reduced mass, and dielectric constant of the surrounding medium using experimental magnetoexciton energies available for monolayer transition metal dichalcogenides (TMDCs). Our approach is based on analytical formulations that allow us to calculate the screening length $r_{0}$ and bandgap energy $E_{g}$ directly from the experimental $s$-state exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$. We also establish a relationship between the surrounding dielectric constant $\kappa$ and the exciton reduced mass $\mu$. This relationship simplifies the Schrödinger equation for a magnetoexciton in a TMDC monolayer, transforming it into a one-parameter equation that depends only on a single material parameter, $\mu$. Therefore, the fitting method is easy to implement, as it compares calculated magnetoexciton energies with experimental data to determine the true exciton reduced mass, the best-fit value of $\mu$. By applying this method, we extract material parameters $E_{g}$, $r_{0}$, $\mu$, and $\kappa$ from the magnetoexciton energies of monolayer TMDCs (WSe2, WS2, MoSe2, and MoS2) encapsulated by hexagonal boron nitride (hBN) slabs in various current experiments. Additionally, we provide key material properties such as diamagnetic coefficients and exciton radii, which are calculated using the extracted parameters.
@@ I Introduction
Two-dimensional semiconductors with thickness of atomic scales, such as monolayer transition metal dichalcogenides (TMDCs), have attracted significant attention over the past decade because they enable the fabrication of ultra-thin electronic devices with exceptional performance [1, 2, 3, 4]. High promises of these materials come from their incredible properties, which include bandgap tunability, high optical sensitivity, strong spin-orbit coupling, valley polarization, considerable excitonic binding energy, and the high thermal stability of exciton at room temperature [5, 6, 7, 8, 9, 10]. Therefore, a precise knowledge of the intrinsic optoelectric properties of these TMDCs, like their energy bandgap $E_{g}$, exciton reduced mass $\mu$, screening length $r_{0}$ (related to the two-dimensional (2D) static polarizability $\chi_{2D}$) as well as the effective dielectric constant $\kappa$ of the surrounding medium is essential for advancing the development of TMDC–based devices.
@@
There are various methods available for determining the optoelectric properties of TMDCs. For instance, the exciton reduced mass can be directly determined via effective electron and hole masses extracted from angle-resolved photoemission spectroscopy (ARPES), which can experimentally detect energy versus momentum maps [11, 12, 13, 14]. Meanwhile, the bandgap energy can be probed by time-resolved ARPES [13], scanning tunneling microscopy (STM) [15, 16, 17], or photoluminescence excitation (PLE) spectra [18, 19]. From the theoretical side, ab initio calculations based on density functional theory (DFT), time-dependent DFT, or GW-Bethe-Salpeter-equation (GW-BSE) could give some predictions about the exciton reduced mass, bandgap energy, or even 2D polarizability [20, 21, 22, 23, 24, 25]. Unfortunately, balancing accuracy and cost among these experimental measurements or theoretical calculations is challenging.
@@
Nevertheless, exciton spectroscopy provides other ways to probe intrinsic optoelectric properties of materials, combining theoretical knowledge of exciton spectra and experimental measurements via optical spectroscopies. Recent studies [26, 27, 28, 29] revealed the exciton reduced mass $\mu$ from the optical spectroscopy of magnetoexcitons at a high magnetic field (65 to 91 tesla) in the same manner as people did in 3D perovskite semiconductors [30, 31]; however, such a magnetic field is not theoretically large enough for their approximation (see a discussion in Ref.  [32]). Another promising approach is to use both PLE and two-photon measurements as people did in 3D semiconductors [31]. This approach requires a theoretical model to describe the relationship between the exciton spectra of $s$ and $p$ states, which is currently unavailable due to the nonhydrogenic nature of exciton spectra in TMDC monolayers. Indeed, in this case, unlike 3D semiconductors, the dimensionality reduction causes a screening effect in the Coulomb interaction between electron and hole [21, 33], which makes them coupled by the Rytova-Keldysh potential [34]. Thus, exciton spectra in TMDC monolayers are much more complicated than the well-known 2D hydrogenic model [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
@@
Figure 1: Schematic flowchart of extracting the material parameters (the exciton reduced mass $\mu$, screening length $r_{0}$, surrounding dielectric constant $\kappa$, and bandgap energy $E_{g}$) from experimental exciton energies.
@@
In the spirit of the availability of magnetoexciton spectroscopy for monolayer TMDCs, our group developed a method of probing material properties, $E_{g}$, $\mu$, $r_{0}$, and $\kappa$, from the magnetoexciton spectra [46, 47, 32]. Particularly, in Ref. [32], the exact numerical calculations of exciton energy spectra via the so-called Feranchuk-Komarov (FK) operator method [48, 49] allowed us to examine the sensitivity of magnetoexciton energies on the material parameters from which we established an efficient fitting scheme for probing material properties. Notwithstanding its accuracy, our numerical fitting scheme relied purely on extensive numerical calculations, preventing us from exploring the underlying physics of the method.
@@
Learning from the case of 3D semiconductors [30, 31], we believe that an analytical model for exciton energy spectra at a zero field can simplify our retrieval process and make sense of the retrieved properties. Fortunately, an analytical expression for s-state exciton energy spectra at zero fields was proposed within the regulated perturbation theory [45]. In the present study, such an analytical formula allows us to develop a formulation for bandgap energy, screening length, and surrounding dielectric constant directly from the zero-field experimental exciton energies. These formulas also enormously reduce the numerical fitting scheme for the exciton reduced mass via magnetoexciton spectra. Figure 1 illustrates our proposal to retrieve material parameters from experimental magnetoexciton energies.
@@
The rest of the present paper is as follows. In Section II, we derive formulas for the bandgap energy $E_{g}$, screening length $r_{0}$, and surrounding dielectric constant $\kappa$. The bandgap energy $E_{g}$ and screening length $r_{0}$ can be calculated directly from zero-field s-state exciton energies. Meanwhile, the formula for the surrounding dielectric constant $\kappa$ also contains the exciton reduced mass $\mu$. Section III is dedicated to the numerical fitting scheme for the exciton reduced mass from the magnetoexciton energies. Section IV provides material parameters $E_{g}$, $r_{0}$, $\mu$, and $\kappa$ retrieved from the experimental exciton energies available for monolayer TMDCs by our method and compares them with the available data. Some other parameters, such as the diamagnetic constants and exciton radii calculated from the retrieved fundamental parameters, are also provided in this section. Section V presents our conclusions.
@@
II Analytical formulas for the bandgap energy, screening length, and surrounding dielectric constant
@@
where Ry $=13605.69$ meV is a Rydberg unit of energy; $P_{n}(\xi)={\textrm{exp}}(a_{n}\Delta)/(1+b_{n}\Delta)$ with $\Delta=0.225/\xi-0.525$, $a_{n}=4.3/n-6.6/n^{2}+3.2/n^{3}$, and $b_{n}=5.3/n-7.6/n^{2}+4.0/n^{3}$. Here, the dimensionless parameter $\eta$ is the rescaling factor of Rydbergy energy, while the dimensionless parameter $\xi$ captures the screening effect from dimensionality reduction. These parameters are defined as
$\eta=\frac{\mu}{\kappa^{2}\,m_{e}},\quad\xi=0.479\,\frac{r_{0}}{a_{0}}\frac{\mu}{\kappa^{2}\,m_{e}},$ (2)
where $\mu$ is the exciton reduced mass; $m_{e}$ is the electron mass; $\kappa$ is the dielectric constant of the surrounding medium; $r_{0}$ is a screening length; $a_{0}=0.0529$ nm is the Bohr radius. We note that formula (1) works well only in the range of $0.290.85>\xi>$ 0.29.
@@
where $E_{g}$ is a bandgap energy of the TMDC monolayer, which needs to be retrieved. From Eqs. (1) and (3), we can inversely define the bandgap energy $E_{g}$ and parameters $\eta$, $\xi$ from the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$ since there are three equations versus three unknowns, as shown by the following scheme:
@@
is a function of only the screening parameter $\xi$ and does not depend on the Rydberg rescaling parameter $\eta$. Therefore, the experimental data of $E_{1s}$, $E_{2s}$, and $E_{3s}$ for different materials such as WSe2, WS2, MoSe2, and MoS2 [27, 29, 37, 38, 28, 36] yield us to the experimental quantity $\delta_{exp}={(E_{3s}-E_{2s})}/(E_{2s}-E_{1s})$, from which we can determine the screening parameter $\xi$ by solving the following equation:
@@
Figure 2: Ratio $\delta(\xi)$ by formula (5) as a function of variable $\xi$ (black line) plotted with some experimental data [27, 28, 36] of $\delta_{\text{exp}}=(E_{2s}-E_{1s})/(E_{3s}-E_{2s})$ marked by different symbols. Also solution $\xi(\delta)$ (7) plotted for comparison (red dots).
@@
$\kappa=\sqrt{\frac{\mu}{\eta\,m_{e}}}\,.$ (11)
Equation (11) does not mean the surrounding dielectric constant $\kappa$ is dependent on the exciton reduced mass $\mu$ but serves for our retrieval scheme only.
For this Section, we conclude that the bandgap energy $E_{g}$ and screening length $r_{0}$ can be determined from the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$ by Eqs. (9) and (10), independently of knowing the values of the surrounding dielectric constant $\kappa$ and the exciton reduced mass $\mu$. This conclusion also explains the numerical results of Ly et al. [32], where $r_{0}$ remains unchanged even when varying the values of $\kappa$ and $\mu$. Besides, the relationship (11) allows reducing one more material parameter, saying the surrounding dielectric constant $\kappa$, from the Schrödinger equation. These three formulas, Eqs. (9), (10), and (11), are essential because they remove the need to determine all structural parameters ($\mu$, $\kappa$, $r_{0}$, and $E_{g}$) simultaneously, remaining to achieve only the exciton reduced mass $\mu$.
@@ III Exciton reduced mass from magnetoexciton energies
To determine the exciton reduced mass $\mu$, we suggest a method based on fitting the calculated magnetoexciton energies with the experimental data.
For the fitting method, we need to solve the following Schrödinger equation for the s-state exciton in monolayer TMDCs as
$\left\{-2\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+\frac{1}{8}\frac{\mu^{2}}{m_{e}^{2}}B^{2}(x^{2}+y^{2})\right.\qquad$
@@
$\left.-\frac{2}{\kappa}\int\limits_{0}^{+\infty}\frac{dq}{\sqrt{1+\alpha^{2}q^{2}}}\;\textrm{e}^{-qr}-\varepsilon\right\}\psi(x,y)=0,\quad$ (12)

### v2

@@ Abstract
We propose a straightforward and highly accurate method for extracting material parameters such as screening length, bandgap energy, exciton reduced mass, and the dielectric constant of the surrounding medium from experimental magnetoexciton energies available for monolayer transition metal dichalcogenides (TMDCs). Our approach relies on analytical formulations that allow us to calculate the screening length $r_{0}$ and bandgap energy $E_{g}$ directly from the experimental $s$-state exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$. We also establish a relationship between the surrounding dielectric constant $\kappa$ and the exciton reduced mass $\mu$. This relationship simplifies the Schrödinger equation for a magnetoexciton in a TMDC monolayer, transforming it into a one-parameter equation that depends solely on the single material parameter $\mu$. Furthermore, we develop an analytical formula with high accuracy for magnetoexciton energies as a function of the exciton reduced mass: $E(B,\mu)$. Then, the inverse of this formula allows us to calculate the exciton reduced mass from experimental data on magnetoexciton energies. By applying this method, we extract key material parameters, $E_{g}$, $r_{0}$, $\mu$, and $\kappa$, from the magnetoexciton energies of monolayer TMDCs, including WSe2, WS2, MoSe2, and MoS2, encapsulated by hexagonal boron nitride (hBN) slabs in various current experiments. The material properties we retrieve complement and correct existing experimental and theoretical data. Additionally, we develop an analytical method for calculating diamagnetic coefficients and exciton radii with high accuracy compared to numerical calculations. Based on this method, we provide diamagnetic coefficients and exciton radii computed using the extracted material parameters.
@@ I Introduction
Two-dimensional semiconductors with thickness of atomic scales, such as monolayer transition metal dichalcogenides (TMDCs), have attracted significant attention over the past decade because they enable the fabrication of ultra-thin electronic devices with exceptional performance [1, 2, 3, 4]. High promises of these materials come from their incredible properties, which include bandgap tunability, high optical sensitivity, strong spin-orbit coupling, valley polarization, considerable excitonic binding energy, and the high thermal stability of excitons at room temperature [5, 6, 7, 8, 9, 10]. Therefore, a precise understanding of the intrinsic optoelectric properties of these TMDCs, such as their energy bandgap $E_{g}$, exciton reduced mass $\mu$, screening length $r_{0}$ (related to the two-dimensional (2D) static polarizability $\chi_{2D}$), and the effective dielectric constant $\kappa$ of the surrounding medium, is essential for advancing the development of TMDC–based devices.
@@
Several methods are available for determining the optoelectric properties of TMDCs. For instance, the exciton reduced mass can be directly determined by extracting effective electron and hole masses from angle-resolved photoemission spectroscopy (ARPES), which allows for the experimental detection of energy versus momentum maps [11, 12, 13, 14]. Meanwhile, the bandgap energy can be probed by time-resolved ARPES [13], scanning tunneling microscopy (STM) [15, 16, 17], or photoluminescence excitation (PLE) spectra [18, 19]. From a theoretical perspective, ab initio calculations based on density functional theory (DFT), time-dependent DFT, or the GW-Bethe-Salpeter equation (GW-BSE) can provide predictions for the exciton reduced mass, bandgap energy, or even 2D polarizability [20, 21, 22, 23, 24, 25, 26]. Unfortunately, balancing accuracy and cost among these experimental measurements or theoretical calculations is challenging.
@@
Nevertheless, exciton spectroscopy offers alternative methods to probe the intrinsic optoelectric properties of materials, combining theoretical knowledge of exciton spectra with experimental measurements via optical spectroscopies. Recent studies [27, 28, 29, 30] have revealed the exciton reduced mass $\mu$ from optical spectroscopy of magnetoexcitons at high magnetic fields (65 to 91 tesla), like that people did for 3D perovskite semiconductors [31, 32]. However, such magnetic fields are not theoretically large enough for their approximation (see a discussion in Ref.  [33]). Another promising approach is to utilize both PLE and two-photon measurements, as has been done in 3D semiconductors [32]. This approach requires a theoretical model to describe the relationship between the exciton spectra of $s$ and $p$ states, which is currently unavailable due to the nonhydrogenic nature of exciton spectra in TMDC monolayers. Indeed, in this case, unlike 3D semiconductors, the dimensionality reduction causes a screening effect in the Coulomb interaction between electron and hole [21, 34, 35], which makes them coupled by the Rytova-Keldysh potential [36]. Thus, exciton spectra in TMDC monolayers are much more complicated than the well-known 2D hydrogenic model [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47].
@@
Figure 1: Schematic flowchart of extracting the material parameters (the bandgap energy $E_{g}$, screening length $r_{0}$, exciton reduced mass $\mu$, and surrounding dielectric constant $\kappa$) from experimental exciton energies. Here, for the exciton reduced mass in the last step, we use the analytical formulation (17) and (III) for $\mu$ from the experimental data.
@@
In the spirit of the availability of magnetoexciton spectroscopy for monolayer TMDCs, our group developed a method of probing material properties, $E_{g}$, $\mu$, $r_{0}$, and $\kappa$, from the magnetoexciton spectra [48, 49, 33]. Particularly, in Ref. [33], the exact numerical calculations of exciton energy spectra via the so-called Feranchuk-Komarov (FK) operator method [50, 51] allowed us to examine the sensitivity of magnetoexciton energies on the material parameters from which we established an efficient fitting scheme for probing material properties. Notwithstanding its accuracy, our numerical fitting scheme relied purely on extensive numerical calculations, preventing us from exploring the underlying physics of the method. Besides, due to the high computing resource consumption of the three-parameter fitting procedure, Ref. [33] provides only results for two TMDC monolayers: WSe2 and WS2.
@@
Learning from the case of 3D semiconductors [31, 32], we believe that an analytical model for exciton energy spectra can simplify our retrieval process and provide insight into the retrieved properties. Fortunately, an analytical expression for s-state exciton energy spectra at zero field was proposed within the regulated perturbation theory [47]. In the present study, this analytical formula enables us to derive a formulation for bandgap energy, screening length, and surrounding dielectric constant directly from the zero-field experimental exciton energies. These formulas also allow us to apply a one-parameter fitting procedure for the exciton reduced mass using magnetoexciton spectra, which significantly reduces the computer resources compared to the previous three-parameter fitting scheme. From other hand, with the achieved analytical formulas, the Schrödinger equation for a magnetoexciton remains containing only one parameter, $\mu$. We can solve this equation to obtain an analytical expression for energies as functions of the exciton reduced mass $E(B,\mu)$. This one-variable function for each given magnetic intensity can be inversely transformed to obtain the exciton reduced mass as a function of exciton energies, which is a key factor in determining the exciton reduced mass from experimental magnetoexciton energies using an analytical method, without fitting. Figure 1 illustrates our proposal to retrieve material parameters from experimental magnetoexciton energies.
@@
The rest of this paper is as follows. In Section II, we derive formulas for the bandgap energy $E_{g}$, screening length $r_{0}$, and surrounding dielectric constant $\kappa$. These formulas allow for the direct calculation of the bandgap energy $E_{g}$ and screening length $r_{0}$ from zero-field s-state exciton energies. Meanwhile, the formula for the surrounding dielectric constant $\kappa$ also contains the exciton reduced mass $\mu$. Section III develops an analytical method for the exciton reduced mass $\mu$ from magnetoexciton energies based on formulations derived in this section. Section IV presents the material parameters $E_{g}$, $r_{0}$, $\mu$, and $\kappa$ retrieved from the experimental exciton energies available for monolayer TMDCs using our method, and compares them with the available data. Some discussions are also given. Other parameters, such as the diamagnetic constants and exciton radii calculated from the retrieved fundamental parameters, are also provided in this section. Section V presents our conclusions and outlook.
@@
II Analytical formulas for bandgap energy, screening length, and surrounding dielectric constant
@@
where Ry =13605.69=13605.69 meV is a Rydberg unit of energy; $P_{n}(\xi)={\textrm{exp}}(a_{n}\Delta)/(1+b_{n}\Delta)$ with $\Delta=0.225/\xi-0.525$, $a_{n}=4.3/n-6.6/n^{2}+3.2/n^{3}$, and $b_{n}=5.3/n-7.6/n^{2}+4.0/n^{3}$. Here, the dimensionless parameter $\eta$ is the rescaling factor of Rydberg energy, while the dimensionless parameter $\xi$ captures the screening effect from dimensionality reduction. These parameters are defined as
$\eta={\frac{\mu}{\kappa^{2}}},\quad\xi=0.479\,\frac{r_{0}}{a_{0}}\frac{\mu}{\kappa^{2}},$ (2)
where $\mu=\mu^{*}/m_{e}$ is the dimensionless exciton reduced mass; $\kappa$ is the dielectric constant of the surrounding medium; $r_{0}$ is the screening length; $m_{e}$ is the electron mass; and $a_{0}=0.0529$ nm is the Bohr radius. Here, the exciton reduced mass $\mu^{*}$ is calculated from the effective electron and hole masses ($m_{e}^{*}$, $m_{h}^{*}$) by the definition $\mu^{*}=m_{e}^{*}m_{h}^{*}/(m_{e}^{*}+m_{h}^{*})$. We note that Formula (1) works well in the range of $0.85>\xi>0.29$, which captures a broad class of monolayer TMDCs encapsulated by hBN slabs.
@@
where $E_{g}$ is the bandgap energy of the TMDC monolayer, which needs to be retrieved. Using Eqs. (1) and (3), we can inversely define the bandgap energy $E_{g}$ and parameters $\eta$, $\xi$ from the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$ since there are three equations versus three unknowns, as shown by the following scheme:
@@
is a function of only the screening parameter $\xi$ and does not depend on the Rydberg rescaling parameter $\eta$. Therefore, the experimental data of $E_{1s}$, $E_{2s}$, and $E_{3s}$ for different materials such as WSe2, WS2, MoSe2, and MoS2 [28, 30, 39, 40, 29, 38] yield the experimental quantity $\delta_{exp}={(E_{3s}-E_{2s})}/(E_{2s}-E_{1s})$, from which we can determine the screening parameter $\xi$ by solving the following equation:
@@
Figure 2: Ratio $\delta(\xi)$ by Formula (5) as a function of variable $\xi$ (black line) plotted with some experimental data [28, 30, 39, 38, 40, 29] of $\delta_{\text{exp}}=(E_{2s}-E_{1s})/(E_{3s}-E_{2s})$ marked by different symbols. Solution $\xi(\delta)$ (7) also plotted for comparison (red dots).
@@
$\kappa=\sqrt{\frac{\mu}{\eta}}\,.$ (11)
Equation (11) does not imply that the surrounding dielectric constant $\kappa$ is dependent on the exciton reduced mass $\mu$, but rather serves as a basis for our retrieval scheme only.
For this Section, we conclude that the bandgap energy $E_{g}$ and screening length $r_{0}$ can be determined from the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$ by Eqs. (9) and (10), independently of knowing the values of the surrounding dielectric constant $\kappa$ and the exciton reduced mass $\mu$. This conclusion also explains the numerical results of Ly et al. [33], where $r_{0}$ remains unchanged even when varying the values of $\kappa$ and $\mu$. Besides, the relationship (11) allows reducing one more material parameter, saying the surrounding dielectric constant $\kappa$, from the Schrödinger equation. Consequently, these three formulas (Eqs. (9), (10), and (11)) are essential because they remove the need to determine all structural parameters ($\mu$, $\kappa$, $r_{0}$, and $E_{g}$) simultaneously, leaving only the exciton reduced mass $\mu$.
@@ III Exciton reduced mass from magnetoexciton energies
To determine the exciton reduced mass $\mu$ from the available experimental magnetoexciton energies, Ref. [33] suggests a method based on fitting the calculated magnetoexciton energies with the experimental data while varying three parameters $\mu$, $r_{0}$, and $\kappa$. The analytical formulas (7)-(11) achieved in the present study allow reducing the fitting to a one-parameter problem. It saves computer resources significantly. However, more advanced, we suggest an analytical approach to calculate the exciton reduced mass $\mu$, based on the formula developed in this section for it as a function of magnetoexciton energies.
For this purpose, first, we consider the Schrödinger equation for an s-state exciton in monolayer TMDCs with the presence of a homogeneous magnetic field $B$ perpendicular to the TMDC plane as
$\left\{-\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+\frac{1}{8}B^{2}(x^{2}+y^{2})\right$$\qquad\qquad\quad$
@@
$\left.{-2}\int\limits_{0}^{+\infty}\frac{dq}{\sqrt{1+\alpha^{2}q^{2}}}\;\textrm{e}^{-qr}-\varepsilon\right\}\psi(x,y)=0,\quad$ (12)